# Robust CLIP: Unsupervised Adversarial Fine-Tuning of Vision Embeddings for Robust Large Vision-Language Models

### v1

@@ Abstract
Multi-modal foundation models like OpenFlamingo, LLaVA, and GPT-4 are increasingly used for various real-world tasks. Prior work has shown that these models are highly vulnerable to adversarial attacks on the vision modality. These attacks can be leveraged to spread fake information or defraud users, and thus pose a significant risk, which makes the robustness of large multi-modal foundation models a pressing problem. The CLIP model, or one of its variants, is used as a frozen vision encoder in many vision-language models (VLMs), e.g. LLaVA and OpenFlamingo. We propose an unsupervised adversarial fine-tuning scheme to obtain a robust CLIP vision encoder, which yields robustness on all vision down-stream tasks (VLMs, zero-shot classification) that rely on CLIP. In particular, we show that stealth-attacks on users of VLMs by a malicious third party providing manipulated images are no longer possible once one replaces the original CLIP model with our robust one. No retraining or fine-tuning of the VLM is required. The code and robust models are available on GitHub.
@@ 1 Introduction
Figure 1: (Robust) performance of LLaVA-1.5 on vision-language tasks and zero-shot (robust) classification for different CLIP models as vision encoder: (i) the original CLIP, (ii) TeCoA2: robust CLIP with supervised adversarial fine-tuning (Mao et al., 2023) at $\ell_{\infty}$ radius of $\nicefrac{{2}}{{255}}$, and (iii) FARE2: robust CLIP using our proposed unsupervised adversarial fine-tuning at $\ell_{\infty}$ radius of $\nicefrac{{2}}{{255}}$. The original CLIP is completely non-robust. Our FARE2 model has better clean and robust performance than TeCoA2 on almost all downstream tasks, see Fig. LABEL:fig:teaser-attack for qualitative outputs.
@@
Several recent foundation models are trained to semantically align inputs from different modalities in a joint embedding space. The most relevant example is CLIP (Radford et al., 2021), which learns, via contrastive training, to encode text and images into a feature space where inputs, in either form, capturing similar concepts are mapped to be close to each other. These models show great promise for many downstream tasks, in particular thanks to their very good performance in zero-shot settings: for example, they can encode virtually any class via its textual description, which makes them well-suited for zero-shot image classification. Additionally, CLIP-like models are an essential component of recent large vision language models (VLMs): in fact, OpenFlamingo (Awadalla et al., 2023) and LLaVA (Liu et al., 2023b, a) are built connecting the frozen vision encoder of the original CLIP with a large language model (MPT (MosaicML, 2023) and Vicuna (Chiang et al., 2023) respectively). These VLMs exhibit excellent zero-shot generalization capabilities, e.g. in image captioning, visual question answering (VQA) and classification from text prompts.
@@
Given the flexibility and effectiveness of such large foundation models, in particular vision-language models, it is foreseeable that they will be used in the near future in many real-world applications. This likely large scale deployment raises questions about the safety and alignment of these systems, and how to prevent the abuse of their abilities and weaknesses by malicious actors. Therefore it becomes extremely important to test and improve the robustness of these models. Recent works (Zhao et al., 2023; Zou et al., 2023) have shown that VLMs are highly vulnerable to adversarial attacks on either text or image inputs. In particular, the vision modality is argued to be the easier one to fool (Carlini et al., 2023): even commercial VLMs like BARD could be attacked successfully with large perturbations (Dong et al., 2023b). Moreover, Schlarmann & Hein (2023) show that imperceptible changes of an image can be used for targeted attacks on VLMs. This can be used by malicious third parties by spreading such images on the web for defrauding users or spreading misinformation on a massive scale.
@@
In this paper, we tackle the vulnerability of the vision modality of VLMs as well as generic adversarial robustness of zero-shot classification using CLIP. To this end, we propose FARE (Fine-tuning for Adversarially Robust Embeddings), an unsupervised fine-tuning scheme for the vision embedding of CLIP to make it robust to adversarial perturbations while also preserving the features of the original CLIP model as much as possible. In this way, we simultaneously achieve two objectives: (i) we can readily replace the original CLIP with our robust CLIP in all down-stream tasks without retraining or fine-tuning since the features on clean inputs are (approximately) preserved. (ii) all down-stream tasks, e.g. zero-shot classification or zero-shot tasks of VLMs, become robust to attacks on the vision modality (see an example in Fig. LABEL:fig:teaser-attack).
@@
The only existing method, TeCoA (Mao et al., 2023), for a robust CLIP vision encoder performs supervised adversarial fine-tuning (using ImageNet) on the zero-shot classifier derived from CLIP (see Sec. 3.2). However, the resulting fine-tuned CLIP model shows significant degradation of zero-shot classification accuracy on datasets different from ImageNet, and on integration into VLMs is detrimental to their performance. In extensive experiments we show that FARE-CLIP preserves much better the clean performance of CLIP on downstream tasks such as zero-shot classification or captioning and visual question answering for the VLMs OpenFlamingo and LLaVA, while having better robustness to $\ell_{\infty}$-bounded attacks (see summary in Fig. 1). In particular, we show that using our FARE-CLIP makes LLaVA robust against imperceptible targeted attacks, see Fig. LABEL:fig:teaser-attack. FARE also leads to lower hallucination rate of LLaVA, and can better solve chain-of-thoughts tasks compared to TeCoA.
@@ 2 Related Work
Many VLMs such as Flamingo (Alayrac et al., 2022), OpenFlamingo (OF) (Awadalla et al., 2023), Fromage (Koh et al., 2023), Mini-GPT-4 (Zhu et al., 2023), LLaVA (Liu et al., 2023b, a) and more (Laurençon et al., 2023; Li et al., 2023a; Chen et al., 2023) have recently appeared. Most of them use a pre-trained large language model (LLM) as well as a large vision encoder such as CLIP. The vision encoder is frozen during training, and only the interaction e.g. via a projection layer or cross-attention is learnt. We focus our evaluation on OF (Awadalla et al., 2023) and LLaVA-1.5 (Liu et al., 2023a) as they both use the original ViT-L/14 CLIP model as vision encoder, similar to (Chen et al., 2023; Li et al., 2023a), but are based on different LLMs: OF on MPT-7B (MosaicML, 2023) and LLaVA on Vicuna-7B (Chiang et al., 2023), a fine-tuned version of Llama (Touvron et al., 2023).
@@
The vulnerability of machine learning models to adversarial attacks is well known and has been extensively studied (Szegedy et al., 2014; Goodfellow et al., 2015). Most existing attacks are on mono-modal models, especially those working on image data (Croce & Hein, 2020) or text (Jia & Liang, 2017; Ebrahimi et al., 2018; Zou et al., 2023; Shen et al., 2023). Adversarial training (Madry et al., 2018) is the most prominent defense against adversarial examples.
Adversarial robustness of VLMs.
In the realm of large VLMs, multiple works have begun to investigate their vulnerability to adversarial attacks (Qi et al., 2023; Carlini et al., 2023; Schlarmann & Hein, 2023; Shayegani et al., 2023; Zhao et al., 2023; Bagdasaryan et al., 2023; Dong et al., 2023a; Bailey et al., 2023; Gu et al., 2024). In Schlarmann & Hein (2023) it is shown that an attacker can use imperceptible perturbations of input images to force the model to produce exact outputs of attackers choosing. In Carlini et al. (2023); Qi et al. (2023) large radii adversarial attacks are proposed for jailbreaking VLMs. Supervised adversarial fine-tuning of CLIP has been investigated in Mao et al. (2023), which is the baseline for our work.
@@
It has been investigated for SimCLR (Chen et al., 2020) models in (Kim et al., 2020; Jiang et al., 2020; Fan et al., 2021), whose methods are based on a contrastive loss formulation. Gowal et al. (2020) propose a self-supervised adversarial training scheme based on BYOL (Grill et al., 2020). Robust classifiers are obtained by adding linear heads to their model. Zhang et al. (2022) propose a two-stage training procedure for SimCLR, with clean training done in the first stage and cosine similarity based adversarial training in the second. In contrast, our method focuses on CLIP and ensures robustness of down-stream tasks even in a zero-shot setting.
@@ 3.1 Robustness of CLIP as Zero-Shot Classifier
Since in this case the text prompts $t_{k}$, are fixed, an image embedding function $\phi$ defines a classifier $f$ via its logits
@@ 3.2 Supervised Adversarial Fine-Tuning
$\mathop{\phi_{\rm{FT}}}\nolimits=\mathop{\rm arg\,min}\limits_{\phi}\sum_{i=1}^{n}\max_{\left\|z-x_{i}\right\|_{\infty}\leq\varepsilon}L_{\textrm{TeCoA}}\left(y_{i},f(\phi,x_{i})\right),$ (2)
@@
This approach has two main problems. First, adversarial training is done with respect to the fixed set of text embeddings of the classes of ImageNet. This does not take into account the effect on other text embeddings, e.g. of categories which are not part of ImageNet, and thus the fine-tuning can lead to heavy distortions with respect to unseen classes, which explains the high losses in standard performance for other downstream zero-shot classification tasks, see Table 4. Second, the loss uses the cosine similarity, which effectively means that it only cares about the projection of the embedding on the hypersphere: one could multiply each $\phi(x)$ by a different scalar factor $\alpha(x)$ and the cosine similarity would be unaffected. Thus during fine-tuning it can happen that the embedding is changed along the radial direction in an arbitrary fashion. As other downstream tasks of CLIP, e.g. VLMs, use the unnormalized embedding this can again lead to huge performance losses. While for the first problem there is no easy solution, the second problem could be solved by retraining the part of the VLM that connects the vision and language components. However, our approach solves both problems at the same time, so that we can get the benefits of our robust CLIP model and maintain good clean performance on all downstream tasks without the need of fine-tuning or retraining.
Table 1: Robustness of vision-language models with different CLIP-models. (Robust) performance of OpenFlamingo and LLaVA for two image captioning and visual question answering tasks. In the last column we show for each CLIP-model the average w.r.t. respective evaluation metrics, with the increase/decrease relative to the respective TeCoA model, introduced in Mao et al. (2023). Both FARE models improve over respective TeCoA models both in clean and robust performance. FARE2 maintains very high clean performance close to the original CLIP model.
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
$\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$
OF-9B
CLIP
79.7
1.5
1.1
60.1
0.7
0.4
23.8
0.0
0.0
48.5
1.8
0.0
53.0
1.0
0.4
[…]
TeCoA2
73.5
31.6
21.2
49.5
14.1
9.5
16.6
3.5
2.1
46.2
23.5
20.5
46.4
17.9
13.3
FARE2
79.1 34.2
19.5
57.7 16.4
8.9
21.6 4.1
1.9
47.0 24.0
17.2
51.4
$\uparrow$5.0
19.7
$\uparrow$1.8
11.9
$\downarrow$1.4
[…]
TeCoA4
66.9
28.5
21.6
40.9
12.0
10.3
15.4
2.1
1.8
44.8
23.6
21.3
41.9
16.5
13.7
FARE4
74.1 30.9 22.8 51.4 15.7 10.5 18.6 3.4 2.9 46.1
23.6
21.0
47.5
$\uparrow$5.6
18.4
$\uparrow$1.9
14.3
$\uparrow$0.6
LLaVA 1.5-7B
CLIP
115.5
4.0
3.1
77.5
1.6
1.0
37.1
0.5
0.0
74.5
2.9
0.0
76.2
2.25
1.0
[…]
TeCoA2
98.4
44.2
30.3
57.1
23.2
15.3
24.1
12.1
8.8
66.9
33.8
21.8
61.6
28.3
19.0
FARE2
109.9 53.6 31.0 71.1 29.5 17.5 31.9 14.7 9.1 71.7 34.9 23.0 71.1
$\uparrow$9.5
33.2
$\uparrow$4.9
20.1
$\uparrow$1.1
[…]
TeCoA4
88.3
50.9
35.3
48.6
27.9
19.5
20.7
12.6
9.3
63.2
41.0 31.7
55.2
33.1
24.0
FARE4
102.4 57.1 40.9 61.6 31.4 22.8 27.6 15.8 10.9 68.3
40.7
30.5
65.0
$\uparrow$9.8
36.2
$\uparrow$3.1
26.3
$\uparrow$2.3
@@ 3.3 Unsupervised Adversarial Fine-Tuning of the Image Embedding
The CLIP embedding has been trained on 400M image-text pairs on the WIT dataset (Srinivasan et al., 2021) and provides very good zero-shot performance. Moreover, downstream tasks like VLMs have been tuned using this embedding. Therefore, our goal is to make the vision encoder robust to adversarial attacks while preserving its output on clean points so that it retains clean zero-shot performance and does not require re-training or fine-tuning of components of downstream tasks, like VLMs. As discussed in the previous section, the supervised fine-tuning is not suited for this. Instead, we introduce an unsupervised adversarial fine-tuning scheme which is not bound to any specific dataset, and does not rely on the text encoder. In the following we denote with $\mathop{\phi_{\rm{Org}}}\nolimits$ the original CLIP encoder. Given an image $x$, we propose the following embedding loss:
@@
This loss enforces that the features of perturbed points $\phi(z)$ stay close to the unperturbed ones $\mathop{\phi_{\rm{Org}}}\nolimits(x)$ of the original CLIP model. Moreover, as $L_{\textrm{FARE}}$ goes to zero, the embedding given by the fine-tuned model for clean images is the same as the one by the original model, that is $\left\|\phi(x)-\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|^{2}_{2}\rightarrow 0$: this implies that the fine-tuned CLIP vision encoder can be plugged into VLMs without influencing their performance. For a set of images $(x_{i})_{i=1}^{n}$, our proposed fine-tuning scheme consists in optimizing
@@
The inner maximization problem in Eq. (3) of this feature-based variant of adversarial training can be solved by PGD. We call our method Fine-tuning for Adversarially Robust Embeddings (FARE).
While we focus here on CLIP and its downstream tasks, our approach can be applied to any foundation model which has an intermediate embedding layer linking modalities.
The following result shows that preserving the image embedding, that is keeping the $\ell_{2}$-distance between original $\mathop{\phi_{\rm{Org}}}\nolimits$ and finetuned embedding $\mathop{\phi_{\rm{FT}}}\nolimits$ small, also preserves the cosine similarities between image and text embeddings.
@@
$\leq$ $\left(\frac{1}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}}+\frac{1}{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}\right)\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)-\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}.$

### v2

@@ Abstract
Multi-modal foundation models like OpenFlamingo, LLaVA, and GPT-4 are increasingly used for various real-world tasks. Prior work has shown that these models are highly vulnerable to adversarial attacks on the vision modality. These attacks can be leveraged to spread fake information or defraud users, and thus pose a significant risk, which makes the robustness of large multi-modal foundation models a pressing problem. The CLIP model, or one of its variants, is used as a frozen vision encoder in many large vision-language models (LVLMs), e.g. LLaVA and OpenFlamingo. We propose an unsupervised adversarial fine-tuning scheme to obtain a robust CLIP vision encoder, which yields robustness on all vision down-stream tasks (LVLMs, zero-shot classification) that rely on CLIP. In particular, we show that stealth-attacks on users of LVLMs by a malicious third party providing manipulated images are no longer possible once one replaces the original CLIP model with our robust one. No retraining or fine-tuning of the down-stream LVLMs is required. The code and robust models are available on GitHub.
@@ 1 Introduction
Figure 1: (Robust) performance of LLaVA-1.5 on vision-language tasks and zero-shot (robust) classification for different CLIP models as vision encoder: (i) the original CLIP, (ii) TeCoA2: robust CLIP with supervised adversarial fine-tuning (Mao et al., 2023) at $\ell_{\infty}$ radius of $\nicefrac{{2}}{{255}}$, and (iii) FARE2: robust CLIP using our proposed unsupervised adversarial fine-tuning at $\ell_{\infty}$ radius of $\nicefrac{{2}}{{255}}$. The original CLIP is completely non-robust. Our FARE2 model has better clean and robust performance than TeCoA2 on almost all down-stream tasks, see Fig. LABEL:fig:teaser-attack for qualitative outputs.
@@
Several recent foundation models are trained to semantically align inputs from different modalities in a joint embedding space. The most relevant example is CLIP (Radford et al., 2021), which learns, via contrastive training, to encode text and images into a feature space where inputs, in either form, capturing similar concepts are mapped to be close to each other. These models show great promise for many down-stream tasks, in particular thanks to their very good performance in zero-shot settings: for example, they can encode virtually any class via its textual description, which makes them well-suited for zero-shot image classification. Additionally, CLIP-like models are an essential component of recent large vision language models (LVLMs): in fact, OpenFlamingo (Awadalla et al., 2023) and LLaVA (Liu et al., 2023b, a) are built connecting the frozen vision encoder of the original CLIP with a large language model (MPT (MosaicML, 2023) and Vicuna (Chiang et al., 2023) respectively). These LVLMs exhibit excellent zero-shot generalization capabilities, e.g. in image captioning, visual question answering (VQA) and classification from text prompts.
@@
Given the flexibility and effectiveness of such large foundation models, in particular LVLMs, it is foreseeable that they will be used in the near future in many real-world applications. This likely large-scale deployment raises questions on the safety and alignment of these systems, and how to prevent the abuse of their abilities and weaknesses by malicious actors. Therefore it becomes extremely important to test and improve the robustness of these models. Recent works (Zhao et al., 2023; Zou et al., 2023) have shown that LVLMs are highly vulnerable to adversarial attacks on either text or image inputs. In particular, the vision modality is argued to be the easier one to fool (Carlini et al., 2023): even commercial LVLMs like BARD could be attacked successfully with large perturbations (Dong et al., 2023). Moreover, Schlarmann & Hein (2023) show that imperceptible changes of an image can be used for targeted attacks on LVLMs. This allows malicious third parties to spread such images on the web for defrauding users or spreading misinformation on a massive scale.
@@
In this paper, we tackle the vulnerability of the vision modality of LVLMs as well as generic adversarial robustness of zero-shot classification using CLIP. To this end, we propose FARE (Fine-tuning for Adversarially Robust Embeddings), an unsupervised fine-tuning scheme for the vision embedding of CLIP to make it robust to adversarial perturbations while also preserving the features of the original CLIP model as much as possible. In this way, simultaneously two objectives are achieved: (i) we can readily replace the original CLIP with our robust CLIP in all down-stream tasks without retraining or fine-tuning since the features on clean inputs are (approximately) preserved. (ii) all down-stream tasks, e.g. zero-shot classification or zero-shot tasks of LVLMs, become robust to attacks on the vision modality (see an example in Fig. LABEL:fig:teaser-attack).
@@
The only existing method, TeCoA (Mao et al., 2023), for a robust CLIP vision encoder performs supervised adversarial fine-tuning (using ImageNet) on the zero-shot classifier derived from CLIP (see Sec. 3.2). However, the resulting fine-tuned CLIP model shows significant degradation of zero-shot classification accuracy on datasets different from ImageNet, and on integration into LVLMs is detrimental to their performance. In extensive experiments we show that FARE-CLIP preserves much better the clean performance of CLIP on down-stream tasks such as zero-shot classification or when used inside LVLMs like OpenFlamingo or LLaVA, while having better robustness to $\ell_{\infty}$-bounded attacks (see summary in Fig. 1). In particular, we show that using our FARE-CLIP makes LLaVA robust against imperceptible targeted attacks, see Fig. LABEL:fig:teaser-attack. FARE also demonstrates robustness to jailbreak attacks, leads to lower hallucination rate of LLaVA, and can better solve chain-of-thoughts tasks compared to TeCoA.
@@ 2 Related Work
Many LVLMs such as Flamingo (Alayrac et al., 2022), OpenFlamingo (OF) (Awadalla et al., 2023), Fromage (Koh et al., 2023), Mini-GPT-4 (Zhu et al., 2023), LLaVA (Liu et al., 2023b, a) and more (Laurençon et al., 2023; Li et al., 2023a; Chen et al., 2023) have recently appeared. Most of them use a pre-trained large language model (LLM) as well as a large vision encoder such as CLIP. The vision encoder is frozen during training, and only the interaction e.g. via a projection layer or cross-attention is learnt. We focus our evaluation on OF (Awadalla et al., 2023) and LLaVA-1.5 (Liu et al., 2023a) as they both use the original ViT-L/14 CLIP model as vision encoder, similar to Chen et al. (2023); Li et al. (2023a), but are based on different LLMs: OF on MPT-7B (MosaicML, 2023) and LLaVA on Vicuna-7B (Chiang et al., 2023), a fine-tuned version of Llama (Touvron et al., 2023).
@@
The vulnerability of machine learning models to adversarial attacks is well known and has been extensively studied (Szegedy et al., 2014; Goodfellow et al., 2015). Adversarial training (Madry et al., 2018) is the most prominent defense against adversarial examples. Most existing attacks focus on uni-modal models, especially those working on image data (Croce & Hein, 2020) or text (Jia & Liang, 2017; Ebrahimi et al., 2018; Zou et al., 2023; Shen et al., 2023). Ban & Dong (2022) propose adversarial perturbations that transfer from pre-trained to fine-tuned models. Moreover, adversarial attacks and defenses for deep metric learning models have also been investigated (Mao et al., 2019; Zhou & Patel, 2022; Zhou et al., 2024).
Adversarial robustness of LVLMs.
In the realm of large vision-language models, multiple works have begun to investigate their vulnerability to adversarial attacks (Qi et al., 2023; Carlini et al., 2023; Schlarmann & Hein, 2023; Shayegani et al., 2023; Zhao et al., 2023; Bagdasaryan et al., 2023; Dong et al., 2023; Bailey et al., 2023; Gu et al., 2024). In Schlarmann & Hein (2023) it is shown that an attacker can use imperceptible perturbations of input images to force the model to produce exact outputs of their choice. In Carlini et al. (2023) and Qi et al. (2023) visual adversarial attacks that allow jailbreaking of LVLMs are proposed. In contrast to our setting, these attacks grant adversaries a large perturbation-radius. Supervised adversarial fine-tuning of CLIP has been investigated by Mao et al. (2023), which is the baseline for our work.
@@
It has been investigated for SimCLR (Chen et al., 2020) models in (Kim et al., 2020; Jiang et al., 2020; Fan et al., 2021; Luo et al., 2023; Xu et al., 2023), whose methods are based on a contrastive loss formulation. Gowal et al. (2020) propose a self-supervised adversarial training scheme based on BYOL (Grill et al., 2020). Robust classifiers are obtained by adding linear heads to their model. Zhang et al. (2022) propose a two-stage training procedure for SimCLR, with clean training done in the first stage and cosine similarity based adversarial training in the second. In contrast, our method focuses on CLIP and ensures robustness of down-stream tasks even in a zero-shot setting by preserving the original embedding.
@@ 3.1 Robustness of CLIP as Zero-Shot Classifier
Since in this case the text prompts $t_{k}$ are fixed, an image embedding function $\phi$ defines a classifier $f$ via its logits
@@ 3.2 Supervised Adversarial Fine-Tuning
$\mathop{\phi_{\rm{FT}}}\nolimits=\mathop{\rm arg\,min}\limits_{\phi}\sum_{i=1}^{n}\max_{\left\|z-x_{i}\right\|_{\infty}\leq\varepsilon}L_{\textrm{TeCoA}}\left(y_{i},f(\phi,z)\right),$ (2)
@@
This approach has two main problems. First, adversarial training is done with respect to the fixed set of text embeddings of the classes of ImageNet. This does not take into account the effect on other text embeddings, e.g. of categories which are not part of ImageNet, and thus the fine-tuning can lead to heavy distortions with respect to unseen classes, which explains the high losses in standard performance for other down-stream zero-shot classification tasks, see Table 4. Second, the loss uses the cosine similarity, which effectively means that it only cares about the projection of the embedding on the hypersphere: one could multiply each $\phi(x)$ by a different scalar factor $\alpha(x)$ and the cosine similarity would be unaffected. Thus during fine-tuning it can happen that the embedding is changed along the radial direction in an arbitrary fashion. As other down-stream tasks of CLIP, e.g. LVLMs (Alayrac et al., 2022; Liu et al., 2023b; Li et al., 2023a), use the unnormalized embedding this can again lead to huge performance losses. While for the first problem there is no easy solution, the second problem could be solved by retraining the part of the LVLM that connects the vision and language components. However, our approach solves both problems at the same time, so that we can get the benefits of our robust CLIP model and maintain good clean performance on all down-stream tasks without the need of fine-tuning or retraining.
@@ 3.3 Unsupervised Adversarial Fine-Tuning of the Image Embedding
The CLIP embedding has been trained on 400M image-text pairs on the WIT dataset (Srinivasan et al., 2021) and provides very good zero-shot performance. Moreover, down-stream tasks like LVLMs have been tuned using this embedding. Therefore, our goal is to make the vision encoder robust to adversarial attacks while preserving its output on clean points so that it retains clean zero-shot performance and does not require re-training or fine-tuning of components of down-stream tasks, like LVLMs. As discussed in the previous section, the supervised fine-tuning is not suited for this. Instead, we introduce an unsupervised adversarial fine-tuning scheme which is not bound to any specific dataset, and does not rely on the text encoder. In the following we denote with $\mathop{\phi_{\rm{Org}}}\nolimits$ the original CLIP encoder. Given an image $x$, we propose the following embedding loss:
@@
This loss enforces that the features of perturbed points $\phi(z)$ stay close to the unperturbed ones $\mathop{\phi_{\rm{Org}}}\nolimits(x)$ of the original CLIP model. Moreover, as $L_{\textrm{FARE}}$ goes to zero, the embedding given by the fine-tuned model for clean images is the same as the one by the original model, that is $\left\|\phi(x)-\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|^{2}_{2}\rightarrow 0$: this implies that the fine-tuned CLIP vision encoder can be plugged into LVLMs without influencing their performance. For a set of images $(x_{i})_{i=1}^{n}$, our proposed fine-tuning scheme consists in optimizing
@@
The inner maximization problem in Eq. (3) of this feature-based variant of adversarial training can be solved by PGD. We call our proposed method Fine-tuning for Adversarially Robust Embeddings (FARE).
While we focus here on CLIP and its down-stream tasks, our approach can be applied to any foundation model which has an intermediate embedding layer linking modalities.
The following result shows that preserving the image embedding, that is keeping the $\ell_{2}$-distance between original $\mathop{\phi_{\rm{Org}}}\nolimits$ and fine-tuned embedding $\mathop{\phi_{\rm{FT}}}\nolimits$ small, also preserves the cosine similarities between image and text embeddings, thereby maintaining zero-shot classification performance.
@@
$\leq$ $\vspace{-1mm}\min\Big{(}\vspace{-1mm}\frac{2}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}},\frac{2}{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}\vspace{-1mm}\Big{)}\vspace{-1mm}\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)-\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}.$
@@
Table 1: Robustness of large vision-language models with different CLIP-models. (Robust) performance of OpenFlamingo and LLaVA for two image captioning and visual question answering tasks. In the last column we show for each CLIP-model the average w.r.t. respective evaluation metrics, with the increase/decrease relative to the respective TeCoA model, introduced in Mao et al. (2023). Both FARE models improve over respective TeCoA models both in clean and robust performance. FARE2 maintains very high clean performance close to the original CLIP model.
$\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$
OF-9B CLIP 79.7 1.5 1.1 60.1 0.7 0.4 23.8 0.0 0.0 48.5 1.8 0.0 53.0 1.0 0.4
[…]
FARE2 79.1 34.2 19.5 57.7 16.4 8.9 21.6 4.1 1.9 47.0 24.0 17.2 51.4 $\uparrow$5.0 19.7 $\uparrow$1.8 11.9 $\downarrow$1.4
[…]
FARE4 74.1 30.9 22.8 51.4 15.7 10.5 18.6 3.4 2.9 46.1 23.6 21.0 47.5 $\uparrow$5.6 18.4 $\uparrow$1.9 14.3 $\uparrow$0.6
LLaVA 1.5-7B CLIP 115.5 4.0 3.1 77.5 1.6 1.0 37.1 0.5 0.0 74.5 2.9 0.0 76.2 2.25 1.0
[…]
FARE2 109.9 53.6 31.0 71.1 29.5 17.5 31.9 14.7 9.1 71.7 34.9 23.0 71.1 $\uparrow$9.5 33.2 $\uparrow$4.9 20.1 $\uparrow$1.1
[…]
FARE4 102.4 57.1 40.9 61.6 31.4 22.8 27.6 15.8 10.9 68.3 40.7 30.5 65.0 $\uparrow$9.8 36.2 $\uparrow$3.1 26.3 $\uparrow$2.3